# AI Agent Behavioral Science

### v1

@@ 4.3 Summary
5 AI Agent Behavior Adaptation
The adaptation of AI agent behavior has become increasingly important as AI systems are expected to operate in alignment with human behaviors, intentions, and social contexts. To achieve this, we describe an adaptation framework that draws from cognitive and behavioral models of human action, inspired by the Fogg Behavior Model fogg2009behavior , aiming to guide and refine AI agent behavior in a more interpretable and human-aligned manner, as shown in Figure 4.
@@ 5 AI Agent Behavior Adaptation
The Fogg Behavior Model posits that a behavior is the result of the simultaneous presence of three elements: ability, motivation, and trigger. In this framework, ability refers to the individual’s competence or capacity to perform a given action; motivation reflects the internal desires or external incentives that drive the individual to act; and trigger is the external stimulus or signal that initiates the behavior at the right moment. Crucially, the model emphasizes that all three elements must co-occur for a behavior to manifest. For example, even if a person is highly motivated, a lack of ability will prevent action; similarly, a competent individual will not act without a clear and timely trigger.
To transfer this model to machine learning systems, we reinterpret each component within the context of AI agent behavior. We regard ability as the foundational competence acquired through large-scale pre-training, which enables the model to perform a broad range of tasks across various domains. Motivation is conceptualized as the reward signals or environmental feedback provided during training or interaction, which influence the AI system’s preferences and drive it to act autonomously in pursuit of desired outcomes. Trigger is understood as explicit human interventions—typically in the form of natural language prompts or task-specific instructions—that guide and shape the AI’s behavior in a targeted direction.
By embedding the structure of the Fogg model, we aim to construct a behavioral adaptation mechanism in which AI agents not only act based on their learned capabilities, but also dynamically respond to motivational signals and human-issued prompts. This triadic structure allows for more flexible, controllable, and human-aligned AI agent behavior. To operationalize this framework, we further explore a variety of optimization techniques that serve to adjust and refine the agent’s behavior. These include reinforcement learning to encode motivational dynamics, prompt learning to enhance responsiveness to triggers, and behavior fine-tuning methods to maintain alignment between model outputs and human behavioral expectations. This integration of human behavioral theory into AI system design offers a promising direction for creating AI agents that behave in ways that are not only intelligent but also contextually appropriate, interpretable, and aligned with human values. In Table 5, we summarize works on AI agent behavior adaptation based on this Fogg-inspired framework.
@@ 5.2 Motivation: Reinforcement Learning
Recently, the application of reinforcement learning (RL) techniques to optimize AI behaviors, particularly those of LLMs, has attracted significant attention. By leveraging RL, the outputs of LLMs can be fine-tuned based on human preference datasets, thereby enhancing their alignment with user expectations. RL is a fundamental paradigm in machine learning, characterized by an agent interacting with an environment to optimize decision-making through trial and error. RL consists of six key components:
@@ 5.5 Summary
This section introduces a framework for AI agent behavior adaptation inspired by the Fogg Behavior Model, focusing on aligning AI actions with human intentions and social contexts. The framework models AI agent behavior through three components: ability, motivation, and trigger. Here, ability is established via large-scale pre-training, providing foundational competence; motivation is realized through RL and fine-tuning methods that embed reward signals and feedback from environmental or human preferences; and trigger refers to prompt-based human interventions, directing AI agents toward contextually suitable actions.
For ability, modern transformer-based models (BERT, ViT, RT-1, etc.) are used to form a robust behavioral foundation, encoding general-purpose knowledge and decision-making capabilities. Motivation leverages RL optimization methods—like RLHF, DPO, and TDPO, as well as fine-tuning strategies like personal-enhanced datasets, adapter-based fine-tuning to dynamically align model outputs with human preferences. Finally, the trigger aspect utilizes sophisticated prompting strategies (Instruction-only, Zero/Few-Shot methods) to precisely and flexibly initiate behaviors in AI systems, particularly beneficial in multi-agent collaboration scenarios.
@@
Building upon the adaptation framework outlined in this section, several promising avenues emerge for future research aimed at enhancing AI behavior alignment with human intentions.
$\bullet$ Prompt design. Advancing prompt tuning methods to achieve more nuanced and context-sensitive triggering of AI behavior constitutes another promising direction. Current approaches (instruction-only, zero-shot, few-shot) demonstrate effectiveness but remain limited in their capacity to handle ambiguous, incomplete, or conflicting human instructions. Future work may explore sophisticated prompting frameworks, including prompt ensembles, adaptive prompt selection, and context-aware prompt generation techniques, to significantly improve AI’s flexibility and precision in interpreting and executing human intentions.
$\bullet$ Robustness in complex environments. While current methods such as RLHF and DPO provide foundational techniques for aligning AI behavior with human feedback, challenges remain regarding scalability, sample efficiency, and generalization across diverse user populations and task scenarios. Therefore, future research should address methods to enhance robustness in RL algorithms, such as meta-reinforcement learning, model-based RL frameworks, and uncertainty-aware policy optimization methods, enabling stable and effective adaptation in complex, real-world environments.
$\bullet$ Long-term adaptation and continuous learning. Existing adaptation mechanisms primarily focus on short-term interactions or static scenarios, neglecting the dynamic and evolving nature of real-world contexts. Therefore, future research should aim to develop AI systems that continuously adapt their behaviors over extended interactions, leveraging memory-augmented models, incremental learning approaches, and knowledge consolidation techniques to maintain consistency, stability, and effectiveness across prolonged usage periods.
@@ 6 AI Agent Behavioral Science for Responsible AI
Building on previous sections, we now highlight how integrating AI agent behavioral science enriches the realization of responsible AI. Traditional approaches to responsible AI often emphasize static ethical guidelines, compliance requirements, or broad governance principles jobin2019global . However, as AI agents become increasingly autonomous, adaptive, and embedded within complex socio-technical systems, these static approaches are insufficient. AI agent behavioral science provides a dynamic, granular perspective – focusing on how motivation, ability, and triggers shape AI behavior in real time – which enables more targeted and effective responsible AI interventions.
In particular, AI agent behavioral science offers tools to proactively design and adjust AI behaviors, ensuring that ethical principles are embedded not only as abstract goals but as concrete, adaptable behavioral patterns. This approach aligns well with the dynamic nature of modern AI systems, which continuously learn and respond to their environments and users. Therefore, embedding behavioral science concepts is essential to navigate the evolving challenges of responsible AI, from mitigating bias to ensuring safety and trustworthiness.
We focus on the following five key dimensions of responsible AI, each examined through the lens of AI agent behavioral science:
@@
In Figure 6, we depict how the principles of responsible AI intersect with various dimensions of AI agent behaviors, incorporating insights from behavior science theories, including single-agent, multi-agent, and human-agent interaction paradigms. This synthesis highlights how these theories, when combined with adaptation strategies, contribute to achieving the objectives of responsible AI. To provide a comprehensive overview, Table 6 summarizes key designs in the relevant literature, mapping each principle to its corresponding behavior dimension and adaptation methods. In the following sections, we will delve into the research landscape under each of the five principles, examining how these adaptation methods can be effectively applied to achieve responsible AI.
Figure 6: The idea and examples of adapting AI agent behavioral science for the measurement and optimization of five goals of responsible AI.
@@
Principle Ref. Dimension of Behavior Adaptation Key Design
Fairness tao2024cultural Single-agent Trigger Cultural bias evaluation
wang2025large Human-AI interaction Trigger Identity group bias evaluation
glickman2024human Human-AI interaction Trigger AI-human feedback loops
hofmann2024ai Single-agent Trigger Masked deception detection
bai2025explicitly Single-agent Trigger Implicit bias evaluation
liprompting Single-agent Trigger Causal fairness prompting
liu2024lidao Single-agent Trigger Attention-inspired bias intervention
raza2023mitigating Single-agent Trigger Bias-mitigating dialogue system
Safety zhou2024larger Single-agent Trigger Overconfidence bias evaluation
steyvers2025large Human-AI interaction Trigger Deception through detailed explanations
xie2023defending Single-agent Trigger System-mode self-reminder
ide2024assessing Single-agent Trigger Random guesser test
xu2023llm Single-agent Trigger Micro-prompt design
karaman2024porover Single-agent Motivation Safety preference optimization
@@
krishna2024understanding Single-agent Trigger Anti-sycophancy prompt engineering
chen2024grath Single-agent Motivation Formulation of debate protocols
@@
ide2024assessing Single-agent Ability Random guesser test
xiao2024behavioral Single-agent Ability Behavioral bias evaluation
jiang2025investigating Single-agent Ability Hybrid moral reasoning
Slack2023TalkToModel Single-agent Trigger Conversational explainability system
Cabrera2023Improve Human-AI interaction Trigger AI behavior description
Li2024Nudge Human-AI interaction Trigger Nudge-based framework
@@
scheurer2024large Single-agent Trigger Deception under pressure
zheng2025cheating Single-agent Trigger LLMs deceive benchmarks
@@
hong2024dpopt Single-agent Motivation Private offsite prompt tuning
staab2024beyond Single-agent Trigger Sensitive attribute inference
zhao2025does Single-agent Trigger Membership inference attack
@@ 6.1 Fairness
Fairness in responsible AI focuses on ensuring that AI systems treat individuals and groups equitably, without unjust biases based on sensitive attributes such as race, gender, culture, or identity mahoney2020ai . It emphasizes the identification, measurement, and mitigation of both explicit and implicit biases in AI systems to prevent discrimination and promote social justice. For LLMs, fairness entails generating outputs that are culturally sensitive, identity-inclusive, and aligned with social values across diverse user groups.
@@
Measuring fairness in AI models helps reveal hidden biases and informs effective mitigation strategies. Recent work has drawn from human behavioral science, cultural psychology, and sociolinguistics to build more comprehensive evaluation frameworks.
@@
Others explore biases emerging during human-agent interaction. Glickman et al. glickman2024human adopt experimental psychology methods to study how AI influences human judgments. They show that interacting with biased AI outputs can amplify human biases, potentially reinforcing social prejudices through feedback loops.
A different line of work investigates implicit and linguistic biases. Hofmann et al.hofmann2024ai use a “masked deception detection” paradigm to identify racial bias toward dialect speakers without explicitly referencing race—revealing discriminatory tendencies embedded in model behaviors. Bai et al.bai2025explicitly employ word association and decision-making tasks from social psychology to uncover unconscious bias, even in the absence of explicit discriminatory content.
@@
To address these biases, researchers have proposed methods that align model behaviors with fairness principles by integrating theories from causal inference, human cognition, and language communication.
@@
Together, these studies provide complementary perspectives: some uncover different types of bias, while others propose targeted interventions inspired by social theories. The interplay between measurement and optimization reflects a feedback loop—where empirical findings guide mitigation designs, and mitigation outcomes help refine fairness evaluation frameworks.
@@ 6.2 Safety
Safety in responsible AI focuses on ensuring that AI systems operate reliably and predictably, minimizing risks and preventing harm to users and society novelli2023account . This involves designing AI behaviors that adhere to safety standards, prevent unintended consequences, and align with human expectations, thereby fostering trust and reliability in AI technologies.
@@
Measuring the safety of AI systems, particularly LLMs, involves assessing their reliability and alignment with human expectations, leveraging insights from behavioral science on perception and decision-making.
Recent studies underscore the gaps between LLM performance and human perception when evaluating their reliability and predictability. Zhou et al. zhou2024larger investigate how scaled-up LLMs, despite enhanced capabilities, produce less predictable and reliable outputs from a human perspective, often generating plausible yet incorrect responses on complex tasks—errors that go unnoticed due to human overconfidence biases akin to those in cognitive psychology. Similarly, Steyvers et al. steyvers2025large explore the misalignment between human trust in LLM outputs and their actual reliability, finding that detailed explanations can inflate user confidence, a phenomenon resembling the halo effect in behavioral research.
In assessing safety through decision-making and contextual influences, other works reveal additional vulnerabilities. Ide et al. ide2024assessing propose the “Random Guesser Test” to evaluate AI safety in sequential decision-making, showing that sophisticated reinforcement learning algorithms can underperform compared to random choices due to limited exploration, mirroring human risk aversion under uncertainty. Xu et al. xu2023llm demonstrate how subtle prompt modifications, such as adding an emoji, can alter LLM outputs, paralleling the impact of minor contextual cues on human behavior in social psychology. Meanwhile, Motwani et al. motwani2024secret uncover risks of LLMs covertly encoding information, evading detection in ways reminiscent of human deception studies. Together, these approaches emphasize a behavioral lens—focusing on trust, error detection, and contextual sensitivity—for evaluating AI safety.
@@
Optimizing AI safety involves refining model behavior and robustness, often drawing on human self-regulation and learning mechanisms from behavioral science.
Recent studies propose techniques inspired by self-regulation to enhance model adherence to safety standards. Xie et al. xie2023defending introduce a “system-mode self-reminder” method, where ethical prompts reinforce ChatGPT’s compliance with safety norms, echoing psychological self-reminders that foster positive behavior. Krishna et al. krishna2024understanding tackle the issue of sycophantic responses in iterative prompting, which undermine truthfulness, and suggest refined prompting strategies—such as repeating questions or extracting facts—to boost accuracy and calibration, reflecting human self-correction processes.
In balancing safety with utility through feedback and validation, other approaches leverage iterative and social mechanisms. Karaman et al. karaman2024porover use overgenerated training data and preference optimization to reduce overrefusal of benign prompts while preserving safety, akin to human learning via reinforced feedback. Brown-Cohen et al. brown2023scalable develop debate protocols where competing AI models justify their outputs to a human verifier, improving safety through argumentation dynamics similar to social influence in behavioral studies. Chen et al. chen2024grath employ out-of-domain prompts to create training data that enhances truthfulness distinctions, using an iterative optimization process that mirrors human trial-and-error learning. These methods align AI safety improvements with behavioral principles like self-discipline, feedback loops, and social validation, providing practical pathways for enhancement.

### v2

@@ 4.3 Summary
5 Adaptation of AI Agent Behaviors
The preceding sections have synthesized emergent AI agent behaviors across three settings: as individuals, in multi-agent environments, and within human-AI interactions. However, understanding behavior is only part of the challenge; equally important is the ability to shape such behavior toward desired goals, values, and contexts. In this section, we shift focus from behavioral observation to behavioral adaptation. That is, we examine methods for guiding and refining AI agent behavior through both traditional learning paradigms and newer agentic design approaches, spanning optimization, instruction, and interaction-level adaptation.
We draw on the Fogg Behavior Model fogg2009behavior , which explains how human behaviors can be changed with the presence of three factors: ability, motivation, and trigger. In the original framework, ability refers to the individual’s competence or capacity to perform a given action; motivation reflects the internal desires or external incentives that drive the individual to act; and trigger is the external stimulus or signal that initiates the behavior at the right moment. Crucially, the model emphasizes that all three elements must co-occur for a behavior to manifest. For example, even if a person is highly motivated, a lack of ability will prevent action; similarly, a competent individual will not act without a clear and timely trigger. We reinterpret these elements in the context of AI agents:
  • Ability maps to foundational competencies acquired during large-scale pretraining, enabling the agent to perform a wide range of tasks.
  • Motivation corresponds to reward signals or environmental feedback introduced via reinforcement learning or strategic fine-tuning, shaping behavioral preferences.
  • Trigger reflects task-specific prompts or instructions that activate and direct agent behavior in specific contexts.
This triadic framework enables us to categorize existing adaptation techniques by the behavioral levers they target. We summarize representative approaches in Table 5, including pretraining for foundational ability, reinforcement learning for motivational alignment, supervised and instruction fine-tuning for contextual value alignment, and prompt engineering for fine-grained behavioral control at inference time. Figure 4 illustrates how these techniques align with the three-part structure of behavioral adaptation.
@@ 5.2 Motivation: Reinforcement Learning
Recently, the application of reinforcement learning (RL) techniques to optimize AI agent behaviors, particularly those of LLMs, has attracted significant attention. By leveraging RL, the outputs of LLMs can be fine-tuned based on human preference datasets, thereby enhancing their alignment with user expectations. RL is a fundamental paradigm in machine learning, characterized by an agent interacting with an environment to optimize decision-making through trial and error. RL consists of six key components:
@@ 5.5 Summary
In this section, we introduce a framework for AI agent behavior adaptation inspired by the Fogg Behavior Model. Our framework models AI agent behavior through three components: ability, motivation, and trigger. Here, ability is established via large-scale pre-training, providing foundational competence; motivation is realized through RL and fine-tuning methods that embed reward signals and feedback from environmental or human preferences; and trigger refers to prompt-based human interventions, directing AI agents toward contextually suitable actions.
For ability, modern transformer-based models (BERT, ViT, RT-1, etc.) are used to form a robust behavioral foundation, encoding general-purpose knowledge and decision-making capabilities. Motivation leverages RL optimization methods—like RLHF, DPO, and TDPO, as well as fine-tuning strategies like personal-enhanced datasets, adapter-based fine-tuning to dynamically align model outputs with human preferences. Finally, the trigger aspect utilizes sophisticated prompting strategies to precisely and flexibly initiate behaviors in AI systems, particularly beneficial in multi-agent collaboration scenarios.
@@
Building upon the adaptation framework outlined in this section, several promising avenues emerge for future research aimed at enhancing AI agent behavior alignment with human intentions.
Prompt design.
Advancing prompt tuning methods to achieve more nuanced and context-sensitive triggering of AI agent behavior constitutes another promising direction. Current approaches (instruction-only, zero-shot, few-shot) demonstrate effectiveness but remain limited in their capacity to handle ambiguous, incomplete, or conflicting human instructions. Future work may explore sophisticated prompting frameworks, including prompt ensembles, adaptive prompt selection, and context-aware prompt generation techniques, to significantly improve AI agents’ flexibility and precision in interpreting and executing human intentions.
Robustness in complex environments.
While current methods such as RLHF and DPO provide foundational techniques for aligning agent behavior with human feedback, challenges remain regarding scalability, sample efficiency, and generalization across diverse user populations and task scenarios. Therefore, future research should address methods to enhance robustness in RL algorithms, such as meta-reinforcement learning, model-based RL frameworks, and uncertainty-aware policy optimization methods, enabling stable and effective adaptation in complex, real-world environments.
Long-term adaptation and continuous learning.
Existing adaptation mechanisms primarily focus on short-term interactions or static scenarios, neglecting the dynamic and evolving nature of real-world contexts. Therefore, future research should aim to develop AI agents that continuously adapt their behaviors over extended interactions, leveraging memory-augmented models, incremental learning approaches, and knowledge consolidation techniques to maintain consistency, stability, and effectiveness across prolonged usage periods.
@@ 6 AI Agent Behavioral Science for Responsible AI
Having examined how AI agent behave across diverse settings and how these behaviors can be adapted, we now turn to a critical application: the pursuit of responsible AI. This section argues that AI Agent Behavioral Science offers a powerful foundation for advancing responsibility in autonomous systems. Traditional approaches to responsible AI often emphasize static ethical guidelines, compliance checklists, or broad governance principles jobin2019global . While necessary, these tools are increasingly insufficient as AI agents become increasingly autonomous, adaptive, and embedded within complex socio-technical systems. A more behaviorally grounded perspective is needed, i.e., one that addresses not just what agents are designed to do, but how they actually behave in practice. AI agent behavioral science fills this gap by offering tools to proactively design and adjust agent behaviors, ensuring that ethical principles are embedded not only as abstract goals but as concrete, adaptable behavioral patterns.
To illustrate this perspective, we focus on five key pillars of responsible AI, each examined through the lens of AI agent behavioral science (see Figure 6):
@@
In the remainder of this section, we review emerging work at this intersection and highlight how behavioral insights can inform concrete interventions and system designs aligned with responsible AI principles. Table 6 summarizes key designs in the relevant literature, mapping each principle to its corresponding behavior dimension and adaptation methods.
Figure 6: Examples of how AI Agent Behavioral Science informs the measurement and optimization of responsible AI principles.
@@
Principle Ref. Scenario Adaptation Key Design
Fairness tao2024cultural Individual agent Trigger Cultural bias evaluation
wang2025large Human-agent interaction Trigger Identity group bias evaluation
glickman2024human Human-agent interaction Trigger AI-human feedback loops
hofmann2024ai Individual agent Trigger Masked deception detection
bai2025explicitly Individual agent Trigger Implicit bias evaluation
liprompting Individual agent Trigger Causal fairness prompting
liu2024lidao Individual agent Trigger Attention-inspired bias intervention
raza2023mitigating Individual agent Trigger Bias-mitigating dialogue system
Safety zhou2024larger Individual agent Trigger Overconfidence bias evaluation
steyvers2025large Human-agent interaction Trigger Deception through detailed explanations
xie2023defending Individual agent Trigger System-mode self-reminder
ide2024assessing Individual agent Trigger Random guesser test
xu2023llm Individual agent Trigger Micro-prompt design
karaman2024porover Individual agent Motivation Safety preference optimization
@@
krishna2024understanding Individual agent Trigger Anti-sycophancy prompt engineering
chen2024grath Individual agent Motivation Formulation of debate protocols
@@
xiao2024behavioral Individual agent Ability Behavioral bias evaluation
jiang2025investigating Individual agent Ability Hybrid moral reasoning
Slack2023TalkToModel Individual agent Trigger Conversational explainability system
Cabrera2023Improve Human-agent interaction Trigger AI behavior description
Li2024Nudge Human-agent interaction Trigger Nudge-based framework
@@
scheurer2024large Individual agent Trigger Deception under pressure
zheng2025cheating Individual agent Trigger LLMs deceive benchmarks
@@
hong2024dpopt Individual agent Motivation Private offsite prompt tuning
staab2024beyond Individual agent Trigger Sensitive attribute inference
zhao2025does Individual agent Trigger Membership inference attack
@@ 6.1 Fairness
Fairness aims to ensure that AI agents treat individuals and groups equitably, avoiding unjust biases based on sensitive attributes such as race, gender, culture, or identity mahoney2020ai . It emphasizes the identification, measurement, and mitigation of both explicit and implicit biases in AI agents to prevent discrimination and promote social justice. Fairness entails generating outputs that are culturally sensitive, identity-inclusive, and aligned with social values across diverse user groups.
@@
Measuring fairness in AI agents provides a foundation for identifying hidden biases and informing mitigation strategies. Recent research has expanded beyond static benchmarks, incorporating methods from experimental psychology and cultural theory to capture biases manifested in interactive and situational settings.
@@
Others explore biases emerging during human-agent interaction. Glickman et al. glickman2024human adopt experimental psychology methods to study how AI agents influences human judgments. They show that interacting with biased AI outputs can amplify human biases, potentially reinforcing social prejudices through feedback loops.
A third line of work focuses on implicit and linguistic biases. Hofmann et al.hofmann2024ai use a “masked deception detection” paradigm to identify racial bias toward dialect speakers without explicitly referencing race—revealing discriminatory tendencies embedded in model behaviors. Bai et al.bai2025explicitly employ word association and decision-making tasks from social psychology to uncover unconscious bias, even in the absence of explicit discriminatory content.
@@
Improving fairness of AI agents requires techniques that integrate fairness principles into both model reasoning and interaction strategies. Recent methods draw inspiration from causal reasoning, cognitive control, and adaptive communication.
@@ 6.2 Safety
Safety focuses on ensuring that AI agents operate reliably and predictably, minimizing risks and preventing harm to users and society novelli2023account . This involves designing agent behaviors that adhere to safety standards and prevent unintended consequences.
@@
Measuring the safety of AI agents, particularly LLM-based ones, involves assessing their reliability and alignment with human expectations, leveraging insights from behavioral science on perception and decision-making.
One line of research highlights the gap between model performance and human perception of reliability. Zhou et al. zhou2024larger investigate how scaled-up LLMs, despite enhanced capabilities, produce less predictable and reliable outputs from a human perspective, often generating plausible yet incorrect responses on complex tasks—errors that go unnoticed due to human overconfidence biases akin to those in cognitive psychology. Similarly, Steyvers et al. steyvers2025large explore the misalignment between human trust in LLM outputs and their actual reliability, finding that detailed explanations can inflate user confidence, a phenomenon resembling the halo effect in behavioral research.
In assessing safety through decision-making and contextual influences, other works reveal additional vulnerabilities. Ide et al. ide2024assessing propose the “Random Guesser Test” to evaluate AI safety in sequential decision-making, showing that sophisticated RL algorithms may perform worse than random choices due to limited exploration, mirroring human risk aversion under uncertainty. Xu et al. xu2023llm demonstrate how subtle prompt modifications, such as adding an emoji, can significantly alter LLM outputs, echoing findings in social psychology on how contextual cues affects human judgment. Meanwhile, Motwani et al. motwani2024secret uncover risks of LLMs covertly encoding information, evading detection in ways parallel to deception mechanisms in human communication. Together, these studies underscore that AI safety must be evaluated with attention to human-like behavioral tendencies, including fallibility, miscalibration, and context dependence.
@@
Optimizing AI safety involves refining agent behavior and robustness, often drawing inspirations from theories of self-regulation, feedback learning, and social accountability.
Some methods are inspired by internal control mechanisms, especially self-regulation. Xie et al. xie2023defending introduce a “system-mode self-reminder” method, where ethical prompts reinforce ChatGPT’s compliance with safety norms, similar to how self-instruction and reminders promote ethical behavior in humans. Krishna et al. krishna2024understanding tackle the issue of sycophantic responses in iterative prompting, which undermine truthfulness, and suggest refined prompting strategies—such as repeating questions or extracting facts—to boost accuracy and calibration, reflecting human self-correction and metacognition.
Other approaches leverage iterative feedback and social validation. Karaman et al. karaman2024porover use overgenerated training data and preference optimization to reduce overrefusal of benign prompts while preserving safety, akin to human learning via reinforced feedback. Brown-Cohen et al. brown2023scalable develop debate protocols where competing AI models justify their outputs to a human verifier, improving safety through argumentation dynamics similar to social influence in behavioral studies. Chen et al. chen2024grath employ out-of-domain prompts to create training data that enhances truthfulness distinctions, using an iterative optimization process that mirrors human trial-and-error learning. By aligning technical interventions with cognitive and social models of safe behavior, these methods offer a pathway toward safer AI agents.